OpenStudy (anonymous):

solve for x radical x-4 =-1i keep gettin 9 :/

OpenStudy (anonymous):

help?

OpenStudy (saifoo.khan):

Lala to rescue!

OpenStudy (saifoo.khan):

\[x \sqrt{x-4}=-1\]

OpenStudy (saifoo.khan):

like that

OpenStudy (saifoo.khan):

?

OpenStudy (anonymous):

noo lozl solve for x radical x-4=-1

OpenStudy (saifoo.khan):

x-4 = 1 x = 5

OpenStudy (saifoo.khan):

Right

OpenStudy (anonymous):

\[\sqrt{x-4}=-1\]

OpenStudy (anonymous):

now you have to confirm the solutions solve the original question

OpenStudy (anonymous):

yaay u guyz are so smart hahaha

OpenStudy (lalaly):

your awesome kiss69fist23 :D:D

OpenStudy (saifoo.khan):

Agree^

OpenStudy (anonymous):

thing is a radical cannot give a negative result unless there is a complex number under the radical

OpenStudy (anonymous):

teehee yaaay

OpenStudy (anonymous):

seems like you're enjoying this material - good for you !

OpenStudy (lalaly):

im enjoying this thread lol

OpenStudy (anonymous):

math is fun !

OpenStudy (lalaly):

\[\huge{WOHOOOOO}\]

OpenStudy (anonymous):

so did you learn about complex numbers kiss69fist23 ?

OpenStudy (anonymous):

im enjoyin it toooooo yaaaaaaay

OpenStudy (anonymous):

btw, you realize that 69/23 is 3 so you can reduce your username to just kiss3fist

myininaya (myininaya):

no real solution

OpenStudy (anonymous):

exactly. thats why I was asking if complex numbers as answers are possible here

myininaya (myininaya):

\[\sqrt{x-4}=-1\] y=(x-4)^(1/2) is above the x-axis y=-1 is below no intersection

OpenStudy (anonymous):

yep - myininaya, you are right !

OpenStudy (anonymous):

Mam Myninaya and sir fiddle are here to help you ............... they are going right i think hence i am not posting my whole solution since it would be better to learn instead of learning the answers